Circular No. 8721

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006cx, 2006cy, 2006cz Additional supernovae have been discovered on unfiltered CCD images:

2006cx and 2006cz by R. R. Prasad, J. Schwehr, and W. Li (LOSS/KAIT; cf. IAUC 8720), and 2006cy by R. Quimby and P. Mondol (cf. IAUC 8622).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006cx June 8.46 22^{h}35^{m}56.^{s}45 +20^{o}19′17′′.1 17.4 1′′.6 E, 3′′.5 S
2006cy June 9.22 13 08 01.23 +26 06 59.0 17.2 2′′.3 E, 13′′.1 S
2006cz June 14.29 14 48 36.66 − 4 44 30.7 17.4 0′′.8 E, 19′′.8 S

Additional KAIT magnitudes for 2006cx in NGC 7316: 2005 Dec. 6.15 UT, [19.0; 2006 June 9.47, 17.5. SN 2006cx is a type-II supernova (details on CBET 543). SN 2006cy, which was found by subtracting a co-addition of 0.45-m ROTSE-IIIb telescope images taken between 2004 Dec. 15 and 2006 Jan. 10 (limiting mag ∼ 19.3) from the 2006 June images, is a type-IIn supernova (details on CBET 544). Additional magnitudes for 2006cz in MCG –01-38-2: May 24.30, [19.1; June 2.31, [18.7; 15.23, 17.3.

COMET C/2006 L2 (McNAUGHT) R. H. McNaught reports his discovery of comet with a 40′′ coma and a

1′.5 tail in p.a. 135^{o} on CCD images taken with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey (discovery observation tabulated below). Following posting on the ‘NEO Confirmation Page’, J. E. McGaha (Tucson, 0.36-m f/10 Schmidt-Cassegrain reflector) writes that seven stacked 30-s CCD images obtained around June 15.15 UT show a bright, stellar nuclear condensation and a coma of diameter 25′′ elongated in p.a. 140^{o}; subsequent images by McGaha around June 15.199-15.205 showed that the brightness was changing by ∼ 1 mag in only 10 min. G. Hug and D. Tibbets (Eskridge, KS, 0.7-m reflector) report that their images from June 15.1 show the object to be diffuse. A visual observation by A. Hale (Cloudcroft, NM, 0.41-m reflector) on June 15.19 revealed a 0′.7 diffuse coma of total mag 13.5.

2006 UT  α_{2000}  δ_{2000}  Mag.
June 14.52128  14^{h}41^{m}31.^{s}41  −38^{o}13′05′′.3  13.7

The available astrometry, preliminary parabolic orbital elements (T = 2006 Nov. 22.076 TT, ω = 51^{o}.353, Ω = 239^{o}.705, i = 99^{o}.708, equinox 2000.0, q = 1.90990 AU), and an ephemeris appear on MPEC 2006-L63.

2006 June 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green